# Color-superconducting quarkyonic matter

### v1

@@ III.1 General formulation of the model
Similarly to normal QY matter, in CSQY matter nucleons occupy the momentum states in the shell of width $\delta_{a}$ above the quark Fermi momentum. In what follows, $k_{a}$ denotes the maximum momentum of nucleons. The latter is assumed to be the same for neutrons and protons. For simplicity, we also ignore its medium dependence.
Following Ref. McLerran and Reddy (2019), we parametrize the width $\delta_{a}$ of the nucleon shell as a piecewise power function of their highest momentum $k_{a}$, i.e.,
@@
where $\Lambda_{a}$ is a constant parameter of the order of the QCD energy scale $\Lambda$ and $\kappa$ is a positive integer. In Refs. McLerran and Reddy (2019); Zhao and Lattimer (2020), a value of $\kappa=2$ was adopted, albeit with parameter choices different from those used here. In contrast to previous parametrizations of the hadronic shell, we present a new perspective on this topic and explore a different range of $\kappa$ values. At $k_{a}<\Lambda_{a}$, the lowest nucleon momentum is zero and, consequently, the corresponding momentum space is nothing but the Fermi sphere with $k_{a}$ being the Fermi momentum of nucleons. At $k_{a}>\Lambda_{a}$, the low-momentum states are not filled by nucleons, which allows populating these states with quarks. Thus, the CSQY matter onsets when the Fermi momentum of nucleons equals $\Lambda_{a}$. Below, the quantities defined at the onset of CSQY matter are labeled with the superscript index $\rm onset$”. For example, the baryon density of the CSQY matter onset is denoted as $n_{B}^{\rm onset}$. This allows us to fix
@@
At high densities, the asymptotic freedom of QCD unbinds quarks from nucleons Gross and Wilczek (1973); Politzer (1973), which leads to the vanishing of nucleon density. Considering that at high maximum momenta of nucleons $k_{a}$, the nucleon shell in the momentum space gets thinner with a width vanishing $\propto 1/k_{a}^{\kappa}$, we conclude that at high densities $n_{a}\simeq\Lambda_{a}^{\kappa+1}k_{a}^{2-\kappa}/\pi^{2}$ in the leading order. At $\kappa=2$ used in Refs. McLerran and Reddy (2019); Zhao and Lattimer (2020) the nucleon densities saturate to the constant values $\Lambda_{a}^{3}/\pi^{2}$, while the nucleon fractions vanish $\propto 1/n_{B}$. Requiring $n_{a}$ to vanish at high densities in agreement with the asymptotic freedom of QCD, in this work, we consider $\kappa\geq 3$.
@@
Below we also consider the speed of sound, which is defined as $c_{S}^{2}=dp/d\varepsilon$, and the dimensionless interaction measure $\delta=1/3-p/\varepsilon$.
In this work, we use the minimal parameterization of the isoscalar nuclear potentials $U_{\rm is}$, which is consistent with the properties of the nuclear matter ground state, i.e., at the nuclear saturation density $n_{0}$ and vanishing isospin asymmetry. Particularly, the vanishing pressure at $n_{0}$, baryon chemical potential $\mu_{0}=M-E_{B}/A$ expressed through the binding energy per nucleon $E_{B}/A$ and nuclear incompressibility factor $K_{0}=9n_{0}^{2}d^{2}(\varepsilon_{0}/n_{0})/dn_{0}^{2}$. Hereafter, the quantities defined at the nuclear matter ground state are labeled with the subscript index “0”. We also account for the nuclear skewness coefficient $Q_{0}=27n_{0}^{3}d^{3}(\varepsilon_{0}/n_{0})/dn_{0}^{3}$, which allows us to modify the stiffness of the nuclear EoS at high densities. To reproduce the present data from nuclear physics and astrophysical constraints, we follow the strategy of a meta-modeling approach Margueron et al. (2018a, b); Zhang et al. (2018), a semi-agnostic, flexible model based on a Taylor expansion of the energy per particle in terms of the particle number density of nucleons and the isospin asymmetry.
@@
Below, we utilize the nucleon mass taken as the mean value of the masses of neutron and proton from the Review of Particle Physics Workman et al. (2022). The properties of the nuclear matter ground state are fixed according to the analysis of the multi-messenger data, together with the nuclear matter experiments within the meta-modeling framework Mondal and Gulminelli (2023). These properties are summarized in Table 1.
@@
Table 1: Nucleon mass and parameters of the nuclear matter ground state utilized in this work and fixed according to Refs. Workman et al. (2022); Mondal and Gulminelli (2023); Scurto et al. (2024).
To provide agreement with the constraints of subsaturation density of the chiral effective field theory Hebeler et al. (2013); Drischler et al. (2016); Tews et al. (2018), we matched the nuclear EoS above with the Haensel-Zdunik EoS (HZ) for the outer crust and the Negele-Vautherin (NV) EoS for the inner crust Haensel et al. (1989); Negele and Vautherin (1973). The matching is provided utilizing the Maxwell crossing of two EoSs.
@@
The low-momentum behavior of $\Delta_{\bf k}$ ensures that the condition $\Delta_{c{\bf k}=0}=0$ is satisfied, leading to $\mu_{fc}=m_{f}$ at vanishing Fermi momenta of quarks, irrespective of their color. At high momenta $\Delta_{\bf k}$ behaves as the momentum-dependent effective gap in nonlocal versions of chiral quark models with the Gaussian formfactor Blaschke et al. (2023b); Contrera et al. (2022); Carlomagno et al. (2024); Ivanytskyi (2025a). The pairing gap (23) has extremum at $|{\bf k}|=\Lambda/\sqrt{2}$. The gap is defined such that its extremal value is $\Delta$, setting the amplitude of the momentum-dependent pairing gap. Hereafter, we refer to $\Delta$ and $\Delta_{\bf k}/\Delta$ as the pairing gap and the scaled momentum-dependent pairing gap, respectively. Fig. 3 shows behavior of $\Delta_{\bf k}/\Delta$. Compared to the Gaussian case, the parametrization (23) leads to a weaker pairing of the low-momentum quark states with $|{\bf k}|\lesssim\Lambda/3$ but enhances pairing of the high-momentum states above this momentum. At the same time, in the nonperturbative range of densities typical for the NS interiors, the high-momentum states are suppressed by the single-particle distribution function of paired quarks. Therefore, the mentioned effects compensate each other and, at a given value of $\Delta$, the parametrization (23) used in this work corresponds to a quark pairing compatible with the Gaussian case.
@@ III.2 Strong equilibrium and color neutrality of color-superconducting quarkyonic matter
The QY picture of dense matter assumes a coexistence of nucleons and quarks. This implies establishing a dynamical equilibrium with respect to strong decays of nucleons to their constituents and the opposite processes of recombination. The corresponding conditions are given by the requirement of equality of chemical potentials of nucleons and the sum of the chemical potentials of their decay products. Such conditions were first introduced to the framework of normal QY matter by Zhao and Lattimer (2020). The condition of strong equilibrium for CSQY matter can be formulated as follows.
@@
Refer to captionRefer to caption
Figure 4: Pressure of electrically neutral CSQY matter at $\beta$-equilibrium as a function of the energy density obtained considering the onset density $n_{B}^{\rm onset}=2n_{0}$ (left panel) and $n_{B}^{\rm onset}=3n_{0}$ (right panel). The calculations are performed for several values of the exponent $\kappa$ and pairing gap $\Delta$, which are indicated in the legend. The purely hadronic EoS is shown by the black dash-dotted curve in both panels. The filled circles indicate the maximum energy densities reached in the centers of the heaviest NSs modeled with the corresponding EoSs. The shaded areas represent the results of the chiral effective field theory Hebeler et al. (2013), perturbative QCD Kurkela et al. (2010) and the constraints extracted from the binary millisecond pulsar PSR J1614-2230 Demorest et al. (2010) and PSR J0740+6620 Fonseca et al. (2021).
@@ III.3 $\beta$-equilibrium and electric neutrality of color-superconducting quarkyonic matter
The NS matter is electrically neutral and exists in equilibrium with respect to direct and inverse $\beta$-decay processes. To incorporate these features into the developed model of CSQY matter, we follow the minimal scheme described in Ref. Zhao and Lattimer (2020) and introduce electrons with mass $m_{e}=0.511$ MeV Workman et al. (2022). Their Fermi momentum $k_{e}$ is found by requiring a zero value of the total density of electric charge, i.e.,
@@ IV Neutron Stars with Color-Superconducting Quarkyonic cores
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 5: The speed of sound $c_{S}^{2}$ (left column) and dimensionless interaction measure $\delta$ (right column) as a functions of energy density $\varepsilon$ obtained for the EoSs shown in Fig. 4. The filled circles indicate the maximum energy densities reached in the centers of the heaviest NSs modeled with the corresponding EoSs. The shaded areas represent the constraints from Ref. Altiparmak et al. (2022); Marczenko et al. (2023) and perturbative QCD Fraga et al. (2014). The idea behind the shaded regions will be explained in the text.
@@ IV.1 Equation of state
With the model parameters fixed, the EoS of CSQY matter corresponding to the conditions in the NS interior can be constructed as outlined in Sec. III. Fig. 4 shows it in the plane of the pressure versus energy density in comparison to the purely hadronic EoS discussed above. As is seen, the pressure of CSQY matter exceeds the hadronic one right after the onset density and gets smaller than the nuclear one at higher densities. This makes the developed model consistent not only with the low-density constraint obtained within the chiral effective field theory Hebeler et al. (2013), but also with the high-density results of perturbative QCD Kurkela et al. (2010). It is also worth mentioning that smaller onset densities lead to the stiffening of the EoS in the density range typical for interiors of NSs. Increasing the pairing gap $\Delta$ also leads to a higher pressure of CSQY matter, which manifests as a CS induced stiffening of its EoS.
@@
Increasing the exponent $\kappa$ leads to a similar effect. While the CS-induced stiffening of the EoS of CSQY matter is caused by the rearrangement of the quark momentum distribution (see Sec. II for discussion), the impact of $\kappa$ warrants a separate remark. It is clear from Eq. (8) that the higher this exponent, the faster the nucleon momentum shell narrows down with an increasing density. This narrowing happens due to unfilling the low-momentum nucleon states and leads to a higher mean momentum of nucleons, which increases the kinetic energy part of the nucleon pressure and, consequently, stiffens the EoS of CSQY matter. In Refs. McLerran and Reddy (2019); Zhao and Lattimer (2020) this effect was observed due to changing $\Lambda$ at $\kappa=2$.
@@
The conclusion regarding the CS-induced stiffening of the EoS of CSQY matter can also be inferred from Fig. 5, showing the speed of sound and dimensionless interaction measure. As is seen, increasing the pairing gap makes the peak of $c_{S}^{2}$ above the onset of CSQY matter more pronounced. Also, the consequent decrease of this quantity becomes steeper with the growth of $\Delta$. This behavior of $c_{S}^{2}$ is reflected in the dimensionless interaction measure, which undergoes a significant decrease at the onset of CSQY matter. Remarkably, the speed of sound remains superconformal at most of the densities typical for the NS interiors, which is another manifestation of the CS induced stiffening of its EoS.
@@
Fig. 5 shows that for not too large values of $\kappa\leq 5$, the present model of CSQY matter agrees well with the astrophysical constraints on speed of sound and dimensionless interaction measure Marczenko et al. (2023); Altiparmak et al. (2022) up to the densities reached in the centers of the heaviest NSs. These constraints were derived from a sampling of EoSs that are consistent with perturbative QCD and chiral effective field theory, while additionally incorporating observational constraints from NSs. Higher values of the exponent $\kappa$ are disfavored by the constraints and, in the case of a softer hadronic model, can even violate causality (see Appendix A) for details. At the same time, at small values of $\kappa$, i.e., $\kappa<4$, the stiffening of the CSQY matter EoS right above its onset is not very significant, especially at small values of the pairing gap. This can be seen from the lower panel of Fig. 5, which indicates that at $\kappa=5$ and $\Delta=0$ the slope of the speed of sound as a function of energy density remains almost the same in the vicinity of the CSQY matter onset. This allows us to conclude that $\kappa\simeq 5$ is phenomenologically most interesting for modeling NSs.

### v2

@@ III.1 General formulation of the model
Similarly to normal QY matter, in CSQY matter nucleons occupy the momentum states in the shell of width $\delta_{a}$ above the quark Fermi momentum. In principle, this width should be self-consistently determined by relating it to the number of quark states above the Fermi sphere, as has been done in the regime of isospin QCD Ivanytskyi:2025cnn. However, this requires accounting for the quark substructure of nucleons manifested by the dynamically generated medium-dependence of their mass. For the sake of simplicity, in this work we omit this step and, following Ref. McLerran:2018hbz, parametrize the width of the nucleon shell as a piecewise power function of their highest momentum $k_{a}$, i.e.,
@@
where $\Lambda_{a}$ is a constant parameter of the order of the QCD energy scale $\Lambda$ and $\kappa$ is a positive integer. In Refs. McLerran:2018hbz; Zhao:2020dvu, a value of $\kappa=2$ was adopted, albeit with parameter choices different from those used here. In contrast to previous parametrizations of the hadronic shell, we present a new perspective on this topic and explore a different range of $\kappa$ values. At $k_{a}<\Lambda_{a}$, the lowest nucleon momentum is zero and, consequently, the corresponding momentum space is nothing but the Fermi sphere with $k_{a}$ being the Fermi momentum of nucleons. At $k_{a}>\Lambda_{a}$, the low-momentum states are not filled by nucleons, which allows populating these states with quarks. Thus, the CSQY matter onsets when the Fermi momentum of nucleons equals $\Lambda_{a}$. Below, the quantities defined at the onset of CSQY matter are labeled with the superscript index $\rm onset$”. For example, the baryon density of the CSQY matter onset is denoted as $n_{B}^{\rm onset}$. This allows us to fix
@@
At high densities, the asymptotic freedom of QCD unbinds quarks from nucleons Gross:1973id; Politzer:1973fx, which leads to the vanishing of nucleon density. Considering that at high maximum momenta of nucleons $k_{a}$, the nucleon shell in the momentum space gets thinner with a width vanishing $\propto 1/k_{a}^{\kappa}$, we conclude that at high densities $n_{a}\simeq\Lambda_{a}^{\kappa+1}k_{a}^{2-\kappa}/\pi^{2}$ in the leading order. At $\kappa=2$ used in Refs. McLerran:2018hbz; Zhao:2020dvu the nucleon densities saturate to the constant values $\Lambda_{a}^{3}/\pi^{2}$, while the nucleon fractions vanish $\propto 1/n_{B}$. Requiring $n_{a}$ to vanish at high densities in agreement with the asymptotic freedom of QCD, in this work, we consider $\kappa\geq 3$.
@@
Below we also consider the speed of sound, which is defined as $c_{S}^{2}=dp/d\varepsilon$, and the dimensionless interaction measure $\delta=1/3-p/\varepsilon$. It is worth mentioning that determining the chemical potentials and pressure from the corresponding thermodynamic identities ensures the thermodynamic consistency of our approach.
In this work, we use the minimal parameterization of the isoscalar nuclear potentials $U_{\rm is}$, which is consistent with the properties of the nuclear matter ground state, i.e., at the nuclear saturation density $n_{0}$ and vanishing isospin asymmetry. Particularly, the vanishing pressure at $n_{0}$, baryon chemical potential $\mu_{0}=M-E_{B}/A$ expressed through the binding energy per nucleon $E_{B}/A$ and nuclear incompressibility factor $K_{0}=9n_{0}^{2}d^{2}(\varepsilon_{0}/n_{0})/dn_{0}^{2}$. Hereafter, the quantities defined at the nuclear matter ground state are labeled with the subscript index “0”. We also account for the nuclear skewness coefficient $Q_{0}=27n_{0}^{3}d^{3}(\varepsilon_{0}/n_{0})/dn_{0}^{3}$, which allows us to modify the stiffness of the nuclear EoS at high densities. To reproduce the present data from nuclear physics and astrophysical constraints, we follow the strategy of a meta-modeling approach Margueron:2017eqc; Margueron:2017lup; Zhang:2018vrx, a semi-agnostic, flexible model based on a Taylor expansion of the energy per particle in terms of the particle number density of nucleons and the isospin asymmetry.
@@
Below, we utilize the nucleon mass taken as the mean value of the masses of neutron and proton from the Review of Particle Physics ParticleDataGroup:2022pth. The properties of the nuclear matter ground state are fixed according to the analysis of the multi-messenger data, together with the nuclear matter experiments within the meta-modeling framework Mondal:2022cva. The chosen symmetry energy slope agrees with the value extracted from the observational data on NSs using a nonparametric Gaussian process EoS Tang:2025xib. These properties are summarized in Table 1.
@@
Table 1: Nucleon mass and parameters of the nuclear matter ground state utilized in this work and fixed according to Refs. ParticleDataGroup:2022pth; Mondal:2022cva; Scurto:2024ekq.
To provide agreement with the constraints of subsaturation density of the chiral effective field theory Hebeler:2013nza; Drischler:2015eba; Tews:2018kmu, we matched the nuclear EoS above with the Haensel-Zdunik EoS (HZ) for the outer crust and the Negele-Vautherin (NV) EoS for the inner crust 1989A&A...222..353H; Negele:1971vb. The matching is provided utilizing the Maxwell crossing of two EoSs.
@@
The low-momentum behavior of $\Delta_{\bf k}$ ensures that the condition $\Delta_{c{\bf k}=0}=0$ is satisfied, leading to $\mu_{fc}=m_{f}$ at vanishing Fermi momenta of quarks, irrespective of their color. At high momenta $\Delta_{\bf k}$ behaves as the momentum-dependent effective gap in nonlocal versions of chiral quark models with the Gaussian formfactor Blaschke:2022egm; Contrera:2022tqh; Carlomagno:2023nrc; Ivanytskyi:2024zip. The pairing gap (23) has extremum at $|{\bf k}|=\Lambda/\sqrt{2}$. The gap is defined such that its extremal value is $\Delta$, setting the amplitude of the momentum-dependent pairing gap. Hereafter, we refer to $\Delta$ and $\Delta_{\bf k}/\Delta$ as the pairing gap and the scaled momentum-dependent pairing gap, respectively. Fig. 3 shows behavior of $\Delta_{\bf k}/\Delta$. Compared to the Gaussian case, the parametrization (23) leads to a weaker pairing of the low-momentum quark states with $|{\bf k}|\lesssim\Lambda/3$ but enhances pairing of the high-momentum states above this momentum. At the same time, in the nonperturbative range of densities typical for the NS interiors, the high-momentum states are suppressed by the single-particle distribution function of paired quarks. Therefore, the mentioned effects compensate each other and, at a given value of $\Delta$, the parametrization (23) used in this work corresponds to a quark pairing compatible with the Gaussian case.
@@ III.2 Strong equilibrium and color neutrality of color-superconducting quarkyonic matter
The QY picture of dense matter assumes a coexistence of nucleons and quarks. This implies establishing a dynamical equilibrium with respect to strong decays of nucleons to their constituents and the opposite processes of recombination. The corresponding conditions are given by the requirement of equality of chemical potentials of nucleons and the sum of the chemical potentials of their decay products. Such conditions were first introduced to the framework of normal QY matter by Zhao:2020dvu. The condition of strong equilibrium for CSQY matter can be formulated as follows.
@@ III.3 $\beta$-equilibrium and electric neutrality of color-superconducting quarkyonic matter
The NS matter is electrically neutral and exists in equilibrium with respect to direct and inverse $\beta$-decay processes. To incorporate these features into the developed model of CSQY matter, we follow the minimal scheme described in Ref. Zhao:2020dvu and introduce electrons with mass $m_{e}=0.511$ MeV ParticleDataGroup:2022pth. Their Fermi momentum $k_{e}$ is found by requiring a zero value of the total density of electric charge, i.e.,
@@ IV Neutron Stars with Color-Superconducting Quarkyonic cores
It is appropriate to add a brief remark concerning the choice of the onset density of CSQY matter. In general, the QY framework allows any value of $n_{B}^{\rm onset}$. However, stiffening of the NS EoS caused by the onset of QY matter (see e.g. Ref. McLerran:2018hbz for discussion) can lead to superluminal $c_{S}^{2}$ if $n_{B}^{\rm onset}$ is high and speed of sound of hadron matter is high itself. At the same time, having this quantity above the onset density of hyperons, which is about $3n_{0}$ (see e.g. Refs. Yamamoto:2015lwa; Yamamoto:2017wre; Li:2019sxd), requires accounting for these particles in hadronic EoS. To provide reliability of the simple phenomenological parameterization of the used hadronic EoS and to omit the necessity to account for hyperons we restrict the onset density of CSQY matter to the values discussed above. It is necessary to stress that they agree with the results of physics-informed Bayesian analysis of the NS data, suggesting that quarks most probably onset at $2n_{0}$ - $3n_{0}$ Ayriyan:2025rub. Recently, a very similar range of $n_{B}^{\rm onset}$ has been reported within a Bayesian analysis of the observational data on NSs with a nonparametric Gaussian process EoS Tang:2025xib. While that study also suggests higher onset densities of quarks about $5n_{0}$, the analysis does not account for hyperons. Therefore, we follow a more reliable range of $n_{B}^{\rm onset}\simeq 1n_{0}-2n_{0}$, which recently has also been used for constructing an interpolating quark-haron EoS Gao:2025vdc, as a guidance.
[…]
Figure 4: Pressure $p$ of electrically neutral CSQY matter at $\beta$-equilibrium as a function of the energy density $\varepsilon$ at the onset density $n_{B}^{\rm onset}=2n_{0}$ (upper panels) and $n_{B}^{\rm onset}=3n_{0}$ (lower panels) shown without zooming the region right after the onset of CSQY matter (left panels) and with that region zoomed (right panels). The calculations are performed for several values of the exponent $\kappa$ and pairing gap $\Delta$, which are indicated in the legend. The purely hadronic EoS is shown by the black dash-dotted curve in both panels. The filled circles indicate the maximum energy densities reached in the centers of the heaviest NSs modeled with the corresponding EoSs. The shaded areas represent the results of the chiral effective field theory Hebeler:2013nza, perturbative QCD Kurkela:2009gj and the constraints extracted from the binary millisecond pulsar PSR J1614-2230 Demorest:2010bx and PSR J0740+6620 Fonseca:2021wxt.
Figure 5: The speed of sound $c_{S}^{2}$ (left column) and dimensionless interaction measure $\delta$ (right column) as a functions of energy density $\varepsilon$ obtained for the EoSs shown in Fig. 4. The filled circles indicate the maximum energy densities reached in the centers of the heaviest NSs modeled with the corresponding EoSs. The shaded areas represent the constraints from Ref. Altiparmak:2022bke; Marczenko:2022jhl and perturbative QCD Fraga:2013qra. The idea behind the shaded regions will be explained in the text.
@@ IV.1 Equation of state
With the model parameters fixed, the EoS of CSQY matter corresponding to the conditions in the NS interior can be constructed as outlined in Sec. III. Fig. 4 shows it in the plane of the pressure versus energy density in comparison to the purely hadronic EoS discussed above. As is seen, the pressure of CSQY matter exceeds the hadronic one right after the onset density and gets smaller than the nuclear one at higher densities. This makes the developed model consistent not only with the low-density constraint obtained within the chiral effective field theory Hebeler:2013nza, but also with the high-density results of perturbative QCD Kurkela:2009gj. It is also worth mentioning that smaller onset densities lead to the stiffening of the EoS in the density range typical for interiors of NSs. Increasing the pairing gap $\Delta$ also leads to a higher pressure of CSQY matter, which manifests as a CS induced stiffening of its EoS.
@@
Increasing the exponent $\kappa$ leads to a similar effect. While the CS-induced stiffening of the EoS of CSQY matter is caused by the rearrangement of the quark momentum distribution (see Sec. II for discussion), the impact of $\kappa$ warrants a separate remark. It is clear from Eq. (8) that the higher this exponent, the faster the nucleon momentum shell narrows down with an increasing density. This narrowing happens due to unfilling the low-momentum nucleon states and leads to a higher mean momentum of nucleons, which increases the kinetic energy part of the nucleon pressure and, consequently, stiffens the EoS of CSQY matter. In Refs. McLerran:2018hbz; Zhao:2020dvu this effect was observed due to changing $\Lambda$ at $\kappa=2$.
@@
The conclusion regarding the CS-induced stiffening of the EoS of CSQY matter can also be inferred from Fig. 5, showing the speed of sound and dimensionless interaction measure. As is seen, increasing the pairing gap makes the peak of $c_{S}^{2}$ above the onset of CSQY matter more pronounced. Also, the consequent decrease of this quantity becomes steeper with the growth of $\Delta$. Understanding this effect requires explaining how the pairing gap amplitude $\Delta$ affects the condition of strong equilibrium. The latter relates the chemical potentials of nucleons to the chemical potentials of quarks, which include the corresponding Fermi energies $\epsilon_{fc}$ having specific values required for ensuring the condition of strong equilibrium. As is seen from the expression given after Eq. (2), $\epsilon_{fc}$ is a growing function of the quark Fermi momentum $k_{fc}$ and a decreasing function of the pairing gap amplitude. Thus, preserving the value of $\epsilon_{fc}$ needed to maintain the strong equilibrium requires a correlated growth of both $k_{fc}$ and $\Delta$. Consequently, increasing $\Delta$ leads to higher $k_{fc}$. At the same time, at high densities after the onset of CSQY matter, the momentum-dependent pairing gap $\Delta_{c{\bf k}}$ is Gaussian in momentum, which causes its suppression due to the increase of the Fermi momentum. This suppression drives the single particle energy of quarks toward the conformal regime and, consequently, causes a faster convergence of the speed of sound to the conformal value manifested by steepening of $c_{S}^{2}$ after its peak. The described behavior of $c_{S}^{2}$ is reflected in the dimensionless interaction measure, which undergoes a significant decrease at the onset of CSQY matter. Remarkably, the speed of sound remains superconformal at most of the densities typical for the NS interiors, which is another manifestation of the CS induced stiffening of its EoS.
@@
Fig. 5 shows that for not too large values of $\kappa\leq 5$, the present model of CSQY matter agrees well with the astrophysical constraints on speed of sound and dimensionless interaction measure Marczenko:2022jhl; Altiparmak:2022bke up to the densities reached in the centers of the heaviest NSs. These constraints were derived from a sampling of EoSs that are consistent with perturbative QCD and chiral effective field theory, while additionally incorporating observational constraints from NSs. Higher values of the exponent $\kappa$ are disfavored by the constraints and, in the case of a softer hadronic model, can even violate causality (see Appendix A) for details. At the same time, at small values of $\kappa$, i.e., $\kappa<4$, the stiffening of the CSQY matter EoS right above its onset is not very significant, especially at small values of the pairing gap. This can be seen from the lower panel of Fig. 5, which indicates that at $\kappa=5$ and $\Delta=0$ the slope of the speed of sound as a function of energy density remains almost the same in the vicinity of the CSQY matter onset. This allows us to conclude that $\kappa\simeq 5$ is phenomenologically most interesting for modeling NSs.